Tsunamis are fast-moving waves often generated by underwater earthquakes. In the deep ocean their amplitude is barely noticable, but upon reaching shore, they can rise up to the astonishing height of a six-story building. One tsunami, generated off the Aleutian islands in Alaska, had a wavelength of 848 km and traveled a distance of 3860 km in 6.31 h. (a) What was the speed (in m/s) of the wave? For reference, the speed of a 747 jetliner is about 250 m/s. Find the wave's (b) frequency and (c) period.
Answer #1

a) speed = distance / time
speed = 3860*1000 / (6.31*3600)
speed = 3860000 / 22716
speed = 169.9242824 m/s

b) f = c / λ
f = 169.92428 / 848000
f = 2.003824085 * 10^-4 Hz

c) T = 1 / f
T = 1 / (2.003824085 * 10^-4)
T = 4990.458031 s